# Predicting and Understanding College Student Mental Health with Interpretable Machine Learning

### v2

@@ Abstract.
To tackle these challenges, this paper presents I-HOPE, the first Interpretable Hierarchical mOdel for Personalized mEntal health prediction. I-HOPE is a two-stage hierarchical model that connects raw behavioral features to mental health status through five defined behavioral categories as interaction labels. We evaluate I-HOPE on the College Experience Study, the longest longitudinal mobile sensing dataset. This dataset spans five years and captures data from both pre-pandemic periods and the COVID-19 pandemic. I-HOPE achieves a prediction accuracy of 91%, significantly surpassing the 60-70% accuracy of baseline methods. In addition, I-HOPE distills complex patterns into interpretable and individualized insights, enabling the future development of tailored interventions and improving mental health support. The code is available at https://github.com/roycmeghna/I-HOPE.
@@ 1. Introduction
We compare I-HOPE with baseline methods that lack personalized predictions or two-stage feature mappings. I-HOPE achieves an overall prediction accuracy of 91%, significantly surpassing the 60-70% accuracy of baseline methods. I-HOPE simplifies complex patterns into interpretable and personalized insights by mapping behavioral features into interaction labels. For instance, a good 7-hour sleep is linked to better mental health outcomes, while poor sleep correlates with increased anxiety and depression. Walking serves as a stress reliever associated with leisure and relaxation. Phone usage reveals contrasting patterns; interactions at home often indicate positive connections, whereas excessive use in social settings suggests stress or discomfort. The balance between social and personal time is crucial, with shared spaces that promote engagement and personal spaces that allow emotional recharge. Importantly, these behaviors affect individuals differently; some find “Me Time” in working out, while others prefer studying or spending time in their own dorms.
@@ 2. College Experience Study (CES) Dataset
This dataset is especially valuable for our study because it covers a five-year period that includes pre-pandemic years, the COVID-19 pandemic, and the gradual return to normalcy as the pandemic receded. By analyzing behavioral patterns and mental health metrics over different time periods, we can assess and predict how mobile sensing data—collected before, during, and after the COVID-19 pandemic—affects students’ mental health. This analysis provides valuable insights into the pandemic’s impact on mental health and the role of behavior in shaping it.
@@ 10-12 Severe
The EMA survey includes a key mental health metric, Patient Health Questionnaire-4 (PHQ-4) score (Kroenke et al., 2009), which serves as the focal point for our analysis and the mental health outcome we aim to predict in this paper. The PHQ-4 score is a widely recognized screening tool for assessing depressive and anxiety symptoms in clinical settings and epidemiological studies (Löwe et al., 2010). It ranges from 0 to 12, with lower scores indicating better mental health. To enhance interpretability and align with standardization practices (Wicke et al., 2022), we categorize PHQ-4 scores into four levels: normal, mild, moderate, and severe. Each category corresponds to specific PHQ-4 score ranges as shown in Table 1, which allows clearer insight into mental health status. The distribution of data points across the PHQ-4 categories is shown in Figure 1.
Figure 1. Imbalanced data distribution across PHQ-4 categories. PHQ-4 category Normal has the highest count, whereas PHQ-4 category Severe has the lowest.
Features.
The dataset includes 172 features, including smartphone activity, duration spent performing various activities, duration spent at various locations, sleep information, etc. However, not all features directly relate to mental health studies, such as the amplitude of detected audio. Additionally, some features, such as distance traveled and number of locations visited, are ambiguous. After careful screening, we select 35 features for our study. These features, which include contextual information, phone usage, and sleep duration, are chosen based on the following criteria.
  1. (1)
  2. (2)
3. Motivation
@@
Table 3. Interaction labels and their corresponding relevant features.
@@
In this section, we explore the challenges of predicting mental health status by analyzing the CES dataset and identify key insights for effective solutions.
3.1. The Role of Smartphone Usage on Mental Health
To assess whether smartphone usage features can effectively predict mental health outcomes, we train a machine learning model on the CES dataset to predict the PHQ-4 score. Given that the PHQ-4 score has four categories, as illustrated in Table 1, this constitutes a multi-class classification problem. Therefore, we use two smartphone features mentioned above—the number of unlocks and daily phone usage duration—as input features to train a multilayer perceptron (MLP) model, a neural network widely used for classification and prediction (Popescu et al., 2009). The MLP model has 2 input nodes, 3 fully connected hidden layers, and 4 output nodes corresponding to the PHQ-4 categories. The model is trained for 50 epochs until the learning curve stabilizes. We preprocess the data by normalizing the features and ensuring balanced class distributions through oversampling for the PHQ-4 categories. We randomly split the dataset into 80% for training and 20% for testing. Our results reveal that the prediction accuracy is only 28%, which is low. Even after oversampling to achieve balanced class distributions for the PHQ-4 categories, the accuracy rises only to 30%. These results indicate that relying solely on these smartphone usage features is insufficient to predict mental health outcomes accurately.
3.3. Improving Predictions with Additional Features
To improve prediction accuracy, we need to incorporate additional features from the CES dataset. Specifically, we use 35 mannully selected features from the CES dataset related to location information, sleep duration, and activity times (see Column 2 of Table 3). We use the MLP model with 35 input nodes, 3 fully connected hidden layers, and 4 output nodes corresponding to the PHQ-4 categories. The training and test sets are the same as above. The model is trained for 50 epochs until the learning curve stabilizes. This approach improves the prediction accuracy to 60%, which is still not high. We attribute the low accuracy to two reasons: (i) treating the dataset as a whole may have obscured individual differences; (ii) a high correlation among features likely reduce the model’s ability to make accurate predictions. We discuss these below.
Addressing Individual Differences
The CES dataset includes data from 217 individuals, each contributing about 160 data points. Research indicates that individual differences in behavior and preferences can significantly affect predictive modeling outcomes (He et al., 2023; Lengerich et al., 2019). To capture these variations, we develop personalized MLP models by grouping data based on each student’s unique user ID (UID). Each model uses the same 35 features but is trained on the corresponding individual’s data, maintaining the original MLP architecture and train-test split. This personalized approach improves the average prediction accuracy to 70%. However, high feature correlation remains a challenge, introducing redundancy and limiting further accuracy improvements, highlighting the need for additional refinement to address these feature-related issues effectively.
Reducing Feature Correlation
Figure 2. Heatmap for correlations among features.
Highly correlated features introduce redundancy, leading to overfitting and reduced test accuracy while increasing computational complexity. As shown in Figure 2, many features are highly correlated. To address this, we analyze feature importance using random forest, an efficient and interpretable ensemble learning technique (Breiman, 2001). This method builds multiple decision trees and combines their outputs to evaluate feature importance. Figure 3 shows the importance ranking of all 35 features, highlighting the top ones. We select the top 50% of these features to retrain personalized MLP models, adjusting the input dimensions accordingly. However, this results in a drop in prediction accuracy to 65%, indicating that simply excluding certain features may remove critical information necessary for accurate predictions.
Figure 3. Feature importances using random forests.
Key takeaway.
Figure 4. Workflow of the I-HOPE design.
4. I-HOPE
Our analysis has shown that globally selecting features for model training reduces correlation but does not enhance prediction accuracy. This highlights the need for a personalized approach that reduces feature correlation and captures individual-specific patterns, leading to better prediction accuracy and interpretability.
Therefore, we design I-HOPE, the first Interpretable Hierarchical mOdel for Personalized mEntal health prediction. I-HOPE is a two-stage hierarchical model that connects behavioral features to mental health status through an intermediate layer of five interaction labels: Leisure, Me Time, Phone Time, Sleep, and Social Time, which categorize different daily behaviors. This model structure offers two benefits:
  1. (1)
    Dimension Reduction. The interaction labels serve as compact representations of complex behaviors, simplifying the input space while retaining the richness of the data.
  2. (2)
Figure 4 presents the workflow of I-HOPE design. It starts with n raw input features ($n=35$ from the CES dataset), which are mapped to k interaction labels ($k=5$), representing aggregated behavioral categories: Leisure, Me Time, Phone Time, Sleep, and Social Time. To quantify the behavioral insights, we assign a score for each interaction label by introducing a two-step rule-based method. These scores then serve as input to a neural network that predicts one of four mental health categories: Normal, Mild, Moderate, or Severe. In summary, by leveraging the hierarchical model structure, I-HOPE improves both prediction accuracy and model interpretability. For example, I-HOPE can assess whether “walking” or “spending time with friends” in the Leisure category has a greater impact on mental health of an individual. I-HOPE is conducted per-user basis, grouping data by user IDs. Next, we will describe each stage of I-HOPE in detail.
@@ 4.1. Stage 1: Input Features $\rightarrow$ Interaction Labels
In Stage 1, I-HOPE maps raw input features to five interaction labels: Leisure, Me Time, Phone Time, Sleep, and Social Time. These labels group features with statistically significant associations to mental health outcomes (p¡0.05 or p¡0.01) into each label. This approach combines domain knowledge and data-driven insights to capture relevant behavioral dimensions, reduce redundancy, and enhance interpretability while maintaining granularity for personalized analysis. The detailed process is described below.
@@
    Feature selection and initial mapping: We manually select a subset of features from the dataset ($n=35$) to match each interaction label based on their semantic and behavioral relevance. For instance, features like unlock duration and phone conversations belong to Phone Time, while duration of social activities and time spent with friends belong to Social Time. The initial mapping is detailed in Table 3. Notably, some features are assigned to multiple labels, reflecting their multidimensional impact on various behavioral categories. We further employ K-means clustering with k=5 to perform a soft validation, assessing whether five interaction labels are a reasonable choice.
@@
    Rule-based scoring initialization: Since interaction labels are manually created, we need to assign them numerical values for downstreaming quantitative analysis. I-HOPE uses a two-step rule-based scoring method to compute an [InteractionLabel]Score for each label. The first step is to initialize the scores. Specifically, for each feature within a label, we compare its value to the population mean value of that feature; if it exceeds the mean, the score increases by 1. The intuition is that higher feature values suggest stronger relevance to the label and thus contribute more to its score. We illustrate the first step of this method using the Sleep label in Algorithm 1, though the same procedure applies to all interaction labels.
@@
    Rule-based scoring using feature importance: The second step of the rule-based scoring method is to use a random forest model on the features and initialized scores to assess the contribution of individual input features to each interaction label. We choose random forest due to its ability to handle high-dimensional data and capture nonlinear relationships (Breiman, 2001). I-HOPE then uses feature importance generated from random forest to refine the final scores. Specifically, for each feature within a label, we compare its value to the population mean; if it exceeds the mean, the score increases by its corresponding feature importance value. We illustrate the second step of this method using the Sleep label in Algorithm 2, though the same procedure applies to all interaction labels.
@@
4:     SleepScore += FeatImp of Sleep duration
@@
7:     SleepScore += FeatImp of Duration at own dorm
@@
10:     SleepScore += FeatImp of Duration of stillness
@@
13:     SleepScore += FeatImp of Conversation detected at home
@@ 5. Evaluation
In this section, we evaluate the I-HOPE to address the following research questions:
@@
To answer these questions, we used Python tools for data preprocessing (cleaning, transformation, and feature extraction) with Pandas and NumPy. I-HOPE was implemented, trained and tested using TensorFlow, while evaluation metrics such as accuracy and loss are computed with Scikit-learn. We randomly split the dataset into 80% for training and 20% for testing.
The rest of this section is divided into three parts, each addressing a research question with detailed results and insights.
@@ 5.1. Prediction Accuracy
Figure 5. Normalized confusion matrix showing high classification accuracy across all categories from I-HOPE.
We compare I-HOPE against several baseline methods that differ in the training approach (aggregated vs. personalized), feature extraction (all raw features vs. selected features) and model structure (single stage vs. hierarchical). The methods are summarized as follows, all using the previously described neural network architecture with adjusted input dimensions:
  • Baseline 1: A single model trained with all raw features.
  • Baseline 2: Personalized models for each individual using all raw features.
  • Baseline 3: Personalized models for each individual using the top 50% of features based on feature importance from random forest models.
@@
The evaluation results over different metrics are summarized in Table 4. Baseline 1 achieves only 60% accuracy due to feature redundancy and lack of personalization. Baseline 2 improves to 70% but still faces feature correlation and redundancy issues. Baseline 3 reduces feature correlation using feature importance but drops to 65% accuracy by excluding significant features. In contrast, I-HOPE reaches 91% overall prediction accuracy by using interaction labels to capture distinct user behaviors while balancing interpretability and predictive power. We can see that the baseline methods struggle with feature redundancy and adapting to user-specific patterns, while I-HOPE effectively integrates personalized interaction labels and feature importance, showing better precision, recall and hence F1 score too. We further use the confusion matrix to provide a detailed breakdown of I-HOPE’s predictions against the true labels. As shown in Figure 5, I-HOPE accurately classifies a majority (more than 90% in three cases) of data points across all four categories. Overall, the evaluation using multiple metrics highlights the strong predictive performance of I-HOPE.

### v3

@@ Abstract.
To tackle these challenges, this paper presents I-HOPE, the first Interpretable Hierarchical mOdel for Personalized mEntal health prediction. I-HOPE is a two-stage hierarchical model that connects raw behavioral features to mental health status through five defined behavioral categories as interaction labels. We evaluate I-HOPE on the College Experience Study, the longest longitudinal mobile sensing dataset. This dataset spans five years and captures data from both pre-pandemic periods and the COVID-19 pandemic. I-HOPE achieves a prediction accuracy of 91%, significantly surpassing the 60-70% accuracy of baseline methods. In addition, I-HOPE distills complex patterns into interpretable and individualized insights, enabling the future development of tailored interventions and improving mental health support.
@@ 1. Introduction
We compare I-HOPE with baseline methods that lack personalized predictions or two-stage feature mappings. I-HOPE achieves an overall prediction accuracy of 91%, significantly surpassing the 60-70% accuracy of baseline methods. I-HOPE simplifies complex patterns into interpretable and personalized insights by mapping behavioral features into interaction labels. For instance, a good 7-hour sleep is linked to better mental health outcomes, while poor sleep correlates with increased anxiety and depression. Walking serves as a stress reliever associated with leisure and relaxation. Phone usage reveals contrasting patterns; interactions at home often indicate positive connections, whereas excessive use in social settings suggests stress or discomfort. The balance between social and personal time is crucial, with shared spaces that promote engagement and personal spaces that allow emotional recharge. Importantly, these behaviors affect individuals differently; some find “Me Time” in working out, while others prefer studying or spending time in their own dorms. The code is available at https://github.com/roycmeghna/I-HOPE.
@@ 2. College Experience Study (CES) Dataset
This dataset is especially valuable for our study because it covers a five-year period that includes pre-pandemic years, the COVID-19 pandemic, and the gradual return to normalcy as the pandemic receded. By analyzing behavioral patterns and mental health metrics over different time periods, we can assess and predict how mobile sensing data—collected before, during, and after the COVID-19 pandemic—predicts students’ mental health. This analysis provides valuable insights into the pandemic’s impact on mental health and the role of behavior in shaping it.
@@ 10-12 Severe
The EMA survey includes a key mental health metric, Patient Health Questionnaire-4 (PHQ-4) score (Kroenke et al., 2009), which serves as the focal point for our analysis and the mental health outcome we aim to predict in this paper. The PHQ-4 score is a widely recognized screening tool for assessing depressive and anxiety symptoms in clinical settings and epidemiological studies (Löwe et al., 2010). It ranges from 0 to 12, with lower scores indicating better mental health. To enhance interpretability and align with standardization practices (Wicke et al., 2022), we categorize PHQ-4 scores into four levels: normal, mild, moderate, and severe. Each category corresponds to a specific PHQ-4 score range as shown in Table 1, which allows clearer insight into mental health status. The distribution of data points across the PHQ-4 categories is shown in Figure 1.
Figure 1. Data distribution across PHQ-4 categories. Normal has the highest count, whereas Severe has the lowest count.
@@
Features.
The original dataset has 172 features, including smartphone usage, duration spent performing various activities, duration spent at various locations, sleep information, etc. However, not all features directly relate to mental health studies, such as the amplitude of detected audio. Additionally, some features, such as distance traveled and number of locations visited, are ambiguous. After careful screening, we select 45 features for our study. These features, which include contextual information, phone usage, and sleep duration, are chosen based on the following criteria.
  1. (1)
  2. (2)
3. Motivation
In this section, we explore the challenges of predicting mental health status by analyzing the CES dataset and identify key insights for designing effective predictive models.
3.1. The Role of Smartphone Usage on Mental Health
To assess whether smartphone usage features can effectively predict mental health outcomes, we train a machine learning model on the CES dataset to predict the PHQ-4 score. Given that the PHQ-4 score has four categories, as illustrated in Table 1, this constitutes a multi-class classification problem. Therefore, we use two smartphone features mentioned above—the number of unlocks and daily phone usage duration—as input features to train a multilayer perceptron (MLP) model, a neural network widely used for classification and prediction (Popescu et al., 2009). The MLP model has 2 input nodes, 3 fully connected hidden layers, and 4 output nodes corresponding to the PHQ-4 categories. The model is trained for 50 epochs until the learning curve stabilizes. We preprocess the data by normalizing the features and ensuring balanced class distributions through oversampling for the PHQ-4 categories. We randomly split the dataset into 80% for training and 20% for testing. Our results reveal that the prediction accuracy is only 28%, which is low. These results indicate that relying solely on these smartphone usage features is insufficient to predict mental health outcomes accurately.
3.3. Improving Predictions with Additional Features
To improve prediction accuracy, we need to incorporate additional features from the CES dataset. Specifically, we use 45 manually selected features from the CES dataset related to location information, sleep duration, and activity times (see Column 2 of Table 3). We use the MLP model with 45 input nodes, 3 fully connected hidden layers, and 4 output nodes corresponding to the PHQ-4 categories. The training and test sets are the same as above. The model is trained for 50 epochs until the learning curve stabilizes. This approach improves the prediction accuracy to 60%, which is still not high. We attribute the low accuracy to two reasons: (i) treating the dataset as a whole may have obscured individual differences; (ii) a high correlation among features likely reduce the model’s ability to make accurate predictions. We discuss these below.
Addressing Individual Differences
The CES dataset includes data from 217 individuals, each contributing about 160 data points on average. Research indicates that individual differences in behavior and preferences can significantly affect predictive modeling outcomes (He et al., 2023; Lengerich et al., 2019). To capture these variations, we develop personalized MLP model for each student. We focus on students who contribute to at least 160 data points in the dataset, which gives us 121 students. Each model is trained on the corresponding individual’s data with 45 selected features, maintaining the same MLP architecture and train-test split as described above. This personalized approach improves the average prediction accuracy to 70%. This motivates the need for additional feature engineering to improve feature representation and prediction accuracy.
Reducing Feature Correlation
Figure 2. Heatmap for correlations among features.
Highly correlated features introduce redundancy, leading to overfitting and reduced test accuracy while increasing computational complexity. As shown in Figure 2, many features are highly correlated. To address this, we analyze feature importance using random forest, an efficient and interpretable ensemble learning technique (Breiman, 2001). This method builds multiple decision trees and combines their outputs to evaluate feature importance. Figure 3 shows the importance ranking of all 45 features, highlighting the top ones. We select the top 50% of these features to retrain personalized MLP models, adjusting the input dimensions accordingly. However, this results in slight a drop in prediction accuracy by $\sim$5%, indicating that simply excluding certain features may remove critical information necessary for accurate predictions.
Key takeaway.
Figure 3. Feature importances using random forests.
Figure 4. Workflow of the I-HOPE design.
4. I-HOPE
Our analysis has shown that (1) personalized models outperform a global model shared among students; and (2) effective feature engineering efforts are needed to reduce feature correlation and improve prediction accuracy. Therefore, we design I-HOPE, the first Interpretable Hierarchical mOdel for Personalized mEntal health prediction. I-HOPE is a two-stage hierarchical model that connects behavioral features to mental health status through an intermediate layer of five interaction labels: Leisure, Me Time, Phone Time, Sleep, and Social Time, which categorize different daily behaviors. This model structure offers two benefits:
  1. (1)
    Dimension Reduction. The interaction labels serve as compact representations of complex behaviors, simplifying the input space while retaining the richness of the data. The effective feature representation contributes to high prediction accuracy.
  2. (2)
Figure 4 presents the workflow of I-HOPE design. It starts with n input features, which are mapped to k interaction labels ($k=5$), representing aggregated behavioral categories: Leisure, Me Time, Phone Time, Sleep, and Social Time. To quantify the behavioral insights, we assign a score for each interaction label by introducing a two-step rule-based method. These scores then serve as input to a neural network model that predicts one of four mental health categories: Normal, Mild, Moderate, or Severe. In summary, by leveraging the hierarchical model structure, I-HOPE improves both prediction accuracy and model interpretability. For example, I-HOPE can assess whether “walking” or “spending time with friends” in the Leisure category has a greater impact on mental health of an individual. I-HOPE is conducted per-user basis by training and testing on individualized data. Before introducing each stage in detail, we will first describe the feature engineering efforts we have made to prepare the final input dataset for I-HOPE.
Table 3. Interaction labels and their corresponding features.
@@ 4. I-HOPE
4.1. Feature Engineering
In Section 3, we have demonstrated that high quality input features are critical to high prediction accuracy. Based on our previous analysis, we further conduct the following feature engineering procedure to prepare the input dataset for I-HOPE.
In our case, feature engineering is guided by domain relevance, ensuring that derived features carry meaningful interpretations in the context of mental health. For example, rather than using the number of phone unlocks and duration of phone usage as separate features, we introduce a ratio feature: number_of_unlocks / duration_of_usage. A high value of this feature suggests more frequent, brief interactions with the phone, and potentially indicative of restlessness. We apply similar transformation across multiple contextual locations, including home, own dorm, study spaces, others’ dorms, social environments, and overall daily usage. In addition, instead of separately including the number and duration of incoming and outgoing calls, we combine them into a single feature:
Number of incoming calls + outgoing callsCall duration of incoming + outgoing $\frac{\texttt{Number of incoming calls + outgoing calls}}{\texttt{Call duration of incoming + outgoing calls}}.$
This consolidation maintains the behavioral signal while reducing redundancy. Overall, we reduce the 45 features to 35 final features, preserving domain-specific behavioral meaning while improving computational efficiency.
@@ 4.2. Stage 1: Input Features $\rightarrow$ Interaction Labels
In Stage 1, I-HOPE maps input 35 features to five interaction labels: Leisure, Me Time, Phone Time, Sleep, and Social Time. This approach combines domain knowledge and data-driven insights to capture relevant behavioral dimensions, reduce redundancy, and enhance interpretability while maintaining granularity for personalized analysis. The detailed process is described below.
@@
    Initial mapping: We manually map each feature to different each interaction label based on their semantic and behavioral relevance. For instance, features like unlock duration and phone conversations belong to Phone Time, while duration of social activities and time spent with friends belong to Social Time. The initial mapping is detailed in Table 3. Notably, some features are assigned to multiple labels, reflecting their multidimensional impact on various behavioral categories. We further employ K-means clustering with k=5 to perform a soft validation, assessing whether five interaction labels are a reasonable choice.
@@
    Rule-based scoring initialization: Since interaction labels are manually created, we need to assign them numerical values for downstreaming quantitative analysis. I-HOPE uses a two-step rule-based scoring method to compute an [InteractionLabel]Score for each label. The first step is to initialize the scores. Specifically, for each feature within a label, we compare its value to the population mean value of that feature; if it exceeds the mean, the score increases by 1. The intuition is that higher feature values suggest stronger relevance to the label and thus contribute more to its score. We illustrate the first step of this method using the Sleep label in Algorithm 1, although the same procedure applies to all interaction labels.
@@
    Rule-based scoring using feature importance: The second step of the rule-based scoring method is to use a random forest model on the features and initialized scores to assess the contribution of individual input features to each interaction label. We choose random forest due to its ability to handle high-dimensional data and capture nonlinear relationships (Breiman, 2001). I-HOPE then uses feature importance generated from random forest to refine the final scores. For each feature within a label, we compare its value to the population mean; if its value exceeds the mean, we increase the score by its Normalized Weighted Feature Impact (NWFI), which is a weighted measure combining the feature’s random forest-derived importance and its normalized value. The normalization reflects the scaling of feature values, and the weighted indicates the incorporation of random forest feature importance. We illustrate the second step of this method using the Sleep label in Algorithm 2, although the same procedure applies to all interaction labels.
@@
4:     SleepScore += NWFI of Sleep duration
@@
7:     SleepScore += NWFI of Duration at own dorm
@@
10:     SleepScore += NWFI of Duration of stillness
@@
13:     SleepScore += NWFI of Conversation detected at home
@@ 4.3. Stage 2: Interaction Labels $\rightarrow$ PHQ-4 Category
4.4. Implementation
We implement I-HOPE in Python. We use Pandas and NumPy for data preprocessing (cleaning, transformation, and feature extraction) and feature engineering. We implement neural network models using TensorFlow. Evaluation metrics such as accuracy are computed with Scikit-learn.
@@ 5. Evaluation
In this section, we evaluate the I-HOPE to address the following research questions (RQs):
@@
The rest of this section is divided into three parts, each addressing a research question.
@@ 5.1. Prediction Accuracy
Figure 5. Normalized confusion matrix from I-HOPE.
We compare I-HOPE against three baseline methods that differ in the training approach (aggregated vs. personalized), feature extraction (all raw features vs. selected features) and model structure (single stage vs. hierarchical). We evaluate 121 students who controbut at least 160 data points in the dataset. We report results from the best-performing model from 5-fold cross validation for all baselines. The methods are summarized as follows, all using the previously described neural network architecture with adjusted input dimensions:
  • Baseline 1: An aggregated single model trained with selected 45 raw features across 121 students.
  • Baseline 2: Personalized models for each individual using selected 45 raw features.
  • Baseline 3: Personalized models for each individual using the top 50% of 45 features based on feature importance from random forest models.
@@
The evaluation results are summarized in Table 4. Baseline 1 achieves 60% accuracy due to feature redundancy and lack of personalization. Baseline 2 improves to 70% but still faces feature correlation and redundancy issues. Baseline 3 reduces feature correlation using feature importance but drops to 65% accuracy by excluding significant features. In contrast, I-HOPE reaches 91% overall prediction accuracy by using interaction labels to capture distinct user behaviors while balancing interpretability and predictive power. We can see that the baseline methods struggle with feature redundancy and adapting to user-specific patterns, while I-HOPE effectively integrates personalized interaction labels and feature importance. We also breakdown I-HOPE’s predictions against the true labels for each category using a confusion matrix. As shown in Figure 5, I-HOPE achieves high accuracy across all four categories. Overall, the evaluation using multiple metrics highlights the strong predictive performance of I-HOPE.